# Estimating LLM Uncertainty with Logits

### v1

@@ Abstract
In recent years, Large Language Models (LLMs) have seen remarkable advancements and have been extensively integrated across various fields. Despite their progress, LLMs are prone to hallucinations, producing responses that may not be dependable if the models lack sufficient grounding knowledge. To mitigate this issue, methods for estimating uncertainty have been adopted, with a focus on critical tokens as indicators of reliability. Nevertheless, probability-based approaches have shown limitations in assessing token-level reliability due to the erosion of evidence strength information acquired during training. In this paper, we introduce Logits-induced Token Uncertainty (LogU), a novel framework designed to estimate token-specific uncertainty in LLMs in real time, without the need for multiple sampling rounds. By leveraging evidence modeling for the implementation of LogU, we utilize the derived uncertainty measures to steer downstream tasks. Our experimental findings highlight the substantial effectiveness and potential of LogU, marking a significant advancement in addressing the challenge of model hallucinations.
@@ 4.2.1 Settings
LogTokU $85.87\%$ $\Uparrow_{7.6\%}$ $86.22\%$ $\Uparrow_{4.8\%}$ $76.19\%$ $\Uparrow_{6.2\%}$ $83.92\%$ $\Uparrow_{1.6\%}$ $97.55\%$ $\Uparrow_{2.8\%}$ $89.11\%$ $\Uparrow_{5.1\%}$
@@ 7 Conclusion
In this paper, we find that the main reason for the failure of probability-based methods in estimating LLM’s uncertainty is that the information regarding evidence strength is lost during the normalization process, making probability cannot accurately reflect reliability. Therefore, we propose LogTokU to model the evidence for generating the next token in LLMs, a real-time token uncertainty estimation framework. We use LogTokU to guide two downstream tasks of LLMs, dynamic decoding and response reliability estimation, achieving significant performance improvements. These downstream tasks demonstrate the simplicity, effectiveness, and immense potential of LogTokU.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v2

@@ Abstract
In recent years, Large Language Models (LLMs) have seen remarkable advancements and have been extensively integrated across various fields. Despite their progress, LLMs are prone to hallucinations, producing responses that may not be reliable if the models lack sufficient grounding knowledge. To mitigate this issue, methods for estimating uncertainty have been adopted, with a focus on critical tokens as indicators of reliability. However, probability-based approaches have shown limitations in assessing token-level reliability due to the loss of evidence strength information acquired during training. In this paper, we introduce Logits-induced Token Uncertainty (LogU), a novel framework designed to estimate token-specific uncertainty in LLMs in real time, without the need for multiple sampling rounds. By leveraging evidence modeling for the implementation of LogU, we utilize the derived uncertainty measures to steer downstream tasks. Our experimental findings highlight the substantial effectiveness and potential of LogU, marking a significant advance in addressing the challenge of model hallucinations.
@@ 4.2.1 Settings
LogTokU $85.87\%$ $\Uparrow_{6.6\%}$ $86.22\%$ $\Uparrow_{11.4\%}$ $76.19\%$ $\Uparrow_{6.2\%}$ $83.92\%$ $\Uparrow_{1.6\%}$ $97.55\%$ $\Uparrow_{2.8\%}$ $89.11\%$ $\Uparrow_{5.1\%}$
@@ 7 Conclusion
In this paper, we find that the main reason for the failure of probability-based methods in estimating LLM’s uncertainty is that the information regarding evidence strength is lost during the normalization process, making probability cannot accurately reflect reliability. Therefore, we propose LogU to model the evidence for generating the next token in LLMs, a real-time token uncertainty estimation framework. We use LogU to guide two downstream tasks of LLMs, dynamic decoding and response reliability estimation, achieving significant performance improvements. These downstream tasks demonstrate the simplicity, effectiveness, and immense potential of LogU.